OpenStudy (anonymous):

6(-3-x)-2x=14 x=? please help i will fan&medal

OpenStudy (saylilbaby):

x=-4

OpenStudy (johnweldon1993):

We can begin by distributing the 6 into the parenthesis \[\large -18 - 6x - 2x = 14\] Now combine like terms...the -6x - 2x = 8x so \[\large -18 - 8x = 14\] Now just add 18 to both sides of this equation \[\large -8x = 32\] and finally divide both sides by -8 \[\large x = -4\]

OpenStudy (anonymous):

thanks @Saylilbaby

OpenStudy (anonymous):

thanks @johnweldon1993

OpenStudy (anonymous):

Yes its -4

OpenStudy (ageta):

before i tell you try to figure it out?

OpenStudy (anonymous):

i tired & i didnt know it

OpenStudy (ageta):

yes Answer: x=−4

OpenStudy (ageta):

your welcome

OpenStudy (12man):

yes ageta is correct

OpenStudy (ageta):

it is ok

OpenStudy (anonymous):

i have one more 10x+4=12x-16

OpenStudy (anonymous):

@ageta

OpenStudy (anonymous):

@Saylilbaby

OpenStudy (saylilbaby):

yes @emo__gamer101

OpenStudy (anonymous):

10x+4=12x-16

OpenStudy (saylilbaby):

10

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Any more problems?

OpenStudy (anonymous):

@emo__gamer101

OpenStudy (ageta):

Step 1: Subtract 12x from both sides. 10x+4−12x=12x−16−12x −2x+4=−16 Step 2: Subtract 4 from both sides. −2x+4−4=−16−4 −2x=−20 Step 3: Divide both sides by -2. −2x −2 = −20 −2 x=10 Answer: x=10